OpenStudy (anonymous):

Ok here's another one of these but the fraction is throwing me off :( Help please? Graph the line that passes through the point (5,1) and has the slope -1/5 .

OpenStudy (anonymous):

Use point slope form with slope \(m\) and point \((x_0,y_0)\).\[\Huge y-y_0=m(x-x_0)\]

OpenStudy (campbell_st):

the equation is in the point slope form \[y = -\frac{1}{5}x + b\] substitute x = 5 and y = 1 to find the value of b

OpenStudy (anonymous):

@campbell_st - That's not point-slope form, that's standard form. ;-)

OpenStudy (campbell_st):

\[1 = -\frac{1}{5} \times 5 + b\] just evaluate to find b

OpenStudy (campbell_st):

but it is point slope form... the question doesn't say to use the point slope formula

OpenStudy (anonymous):

@campbell_st - But you said point-slope form. What you gave is called standard form. Just a terminology correction. Your method was fine.

OpenStudy (campbell_st):

point slope form is y = mx + b my equation is \[y = - \frac{1}{5}x + 2\]

OpenStudy (anonymous):

ok here is a trick. Think of the slope like this; if the slope is -1/5, the numerator at the top is the amount that it moves UP (if positive+) or DOWN(if negative - ) to the next point the denominator at the bottom is the amount that it moves Left or Right to the next point. So you can trace it to the next point like this

OpenStudy (campbell_st):

actually both methods are correct and both methods are in point gradient form

OpenStudy (campbell_st):

just my method is slightly quicker than yours...

OpenStudy (anonymous):

@campbell_st @yakeyglee @javawarrior But what would a graph look like?

OpenStudy (campbell_st):

to graph the line... pick 3 x values and create a table...  plot the points (-5, 3) (0, 2) and (5, 1) to graph your line